f(x)= a tan' (1/x-4), 0<=x<4
pi/2 ,x=4
b tan'(2/x-4) ,4sin'(7-x) + a*pi/4 , 6<=x<=8
find a n b if f(x) is continous at [0,8]
basically i wanna know how to solve such a prob. whr a function is defind under differnt domains n check for continuity n differntibility

P.S. i dont know how to look for diffrntiablility ie (lhd /rhd) plz hlppppp

Grade:

1 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

We will learn about differentiability , and the various factors affecting differentiability. We have already defined the derivative of a function f(x) at a particular point 'a' and derivative of f(x) in general for the variable x as f(a) and f(x) respectively. The restriction in both the cases is that 'the limit must exist'. 

If

does not exist, then we say that the function is not differentiable.

If the above limit exists, we say the function f(x) is differentiable.

In order to test the differentiability of a function at a point, the right hand derivative and left hand derivatives are introduced as follows:

Right Hand Derivative

Let f be a function of x (y=f(x)). Let a be a point in the domain of f. The RHD of f at a is defined as

where h>0, provided the limit exists.

Left Hand Derivative

The LHD of f at a is defined as

where h>0, provided the limit exists.

In the above definition showiing differentiabilty, substitute a + h = x, then h = x - a as

Rf '(a) can be rewritten as

Similarly, substitute a - h = x. Lf '(a) can be written as

Differentiability at a Point

Diffferentiability at a point 'a' for a function f(x) if

(i) both Rf '(a) and Lf '(a) exists and finite.

(ii) Rf '(a) = Lf '(a)

Consider the function y=|x|. This function is differentiable on (,0) and (0,), but not differentiable at x = 0.

For x>0, we have

Since the limit exists, f(x) is differentiable at x>0. Similarly, we can show that f(x) is differentiable at x<0.

We shall find RHD and LHD of f(x) at x = 0.

= -1

Therefore y = |x| is not differentiable at x = 0.